# https://github.com/yfzhang114/LLaVA-Align Debiasing Large Visual Language Models

### v1

@@ Abstract
LLMs such as GPT-4 [34], PaLM [7], and LLaMA-2 [47] have evolved to handle both text and images. This expansion into the multimodal domain involves further pre-training with image-text pairs or fine-tuning with specialized vision instruction tuning datasets, resulting in the birth of powerful Large Visual Language Models (LVLMs). These models excel at understanding complex visual patterns and translating them into coherent language. However, despite their capabilities, our investigation reveals a notable issue. The content generated by LVLMs is significantly biased towards the underlying LLMs used during pre-training, rather than being influenced by the input images. Even in situations where the image is entirely noisy or absent, LVLMs confidently generate answers, indicating a bias towards the learned language patterns. This observation is illustrated in Fig. 1, where querying an LVLM about the color of a non-existent fox produces confident yet unrealistic answers like brown, gray, or blue. To explore this bias further, we conduct a toy experiment using questions about color, shape, number, and relationships for 80 entities in the MSCOCO dataset. Our experimental baselines include Noise (fully noisy image and question input), None (no image provided), and Zeros/Ones (image with all zero/one values). Fig. 1 displays the top 15 choices and corresponding probabilities of frequently occurring answers across these cases. Remarkably, even when no image is available or questions are about non-existent or meaningless images, current LVLMs tend to produce specific answers. This indicates a substantial bias originating from the pretraining on LLMs. Such biases bring persistent challenges, particularly the issue of hallucination [13, 26, 26, 43], pose a significant concern, impacting the reliability and applicability of LVLMs. In this paper, we also aim to uncover the underlying reasons behind such a phenomenon, and the results will provide insights into the design of current LVLMs and their training strategies.
@@
In this paper, we present two training-free strategies to mitigate generation biases associated with underlying LLMs. Our first approach introduces a “calibration” step to adjust the output distribution. Specifically, we assess the model’s inclination toward specific answers by inputting a dummy test with no image, as illustrated in Fig. 1, where we replace the image with ’None’ or meaningless tensors. We then determine the calibration parameters to ensure that the image-free input produces uniform scores for each answer without yielding overly confident outputs. This calibration procedure establishes suitable parameters without requiring additional training data, referred to as the ‘‘Post-Hoc debias’’ approach, proving intuitive for classification tasks111In binary classification settings like POPE, we aim for the input with a meaningless image to have ’yes’ and ’no’ with an equal probability of .. However, in open-ended generation, calibrating each token’s logits uniformly, even without image input, is unreasonable due to the inherent correlation between text tokens. In such cases, we extend the Post-Hoc debias method to its ‘‘Debias sampling’’ counterpart. Inspired by contrastive decoding methods [25, 21], we compute the difference between generation token log-probabilities for one with a correct image and another with a meaningless vision input. Both the Post-Hoc debias and debias sampling methods serve as regularization techniques, reducing the generation results’ dependency on pure text or meaningless image input.
@@
Moreover, we have identified a significant level of instability in the performance of current LVLMs across various generative configurations. Our primary hypothesis is that existing evaluations are predominantly based on a default decoding setting, limiting the exploration of the model’s full capabilities. As extensively illustrated in Fig. 2, diverse choices in generative configurations yield substantially different performance outcomes. Notably, adjusting the generation temperature can lead to a notable increase in the average F1 score of POPE-MSCOCO, shifting from 76 to approximately 84.04. This improvement surpasses the previous results achieved by the 13B model and the RLHF-tuned model by a considerable margin, indicating that the current estimation of LVLM capability may be underestimated. Furthermore, our investigation reveals that different models exhibit preferences for distinct generation configurations, highlighting the instability in the current evaluation of LVLMs. This raises concerns about the fairness of evaluations, particularly since they often rely on default generation configurations or selectively choose the best configuration for the proposed model. To maximize the full potential of existing LVLMs, we systematically investigate six LVLM models across four benchmarks, meticulously searching their generative configurations to identify optimal settings. Our results significantly outperform previously reported ones by a substantial margin, underscoring the importance of selecting the best decoding configuration for each model.
@@ 2 Method
Background of Large Vision-Language Models. A LVLM with parameters , conditioned on a query text , a vision input , parametrizes a probability distribution over response sequences . Each response token is sampled from a token probability distribution, namely , where denotes the logits of . Usually, an LLM is first pre-trained on a large, unlabeled text dataset, then an LVLM utilizes the LLM as the text encoder and incorporates vision representations in different ways [3, 23, 9, 12]. Related work about LVLMs and Hallucination in LVLMs can be found in Appendix 0.A.
@@ 2.1 Exploring the Impact of LLM Biases on LVLMs
To assess the influence of LLM biases on LVLMs, we conduct toy experiments with five target questions detailed in Tab. 4. For each entity of the MSCOCO dataset (appendix Tab. 3), we design corresponding prompts and pair them with different types of vision inputs, including unrelated vision input such as fully noisy images (Noise) or entirely black or white images (Zero/One), or replace vision input with pure text, where we either remove all vision tokens (None) or replace them with a meaningless placeholder like </unk> (Unk). As illustrated in Fig. 1 and in the appendix Fig. 13, LVLMs exhibit biases towards specific answers, such as associating “white” with color-related questions, “round” with shape-related questions, and “many” with number-based questions. We further analyze attention mechanisms to explain why models heavily rely on LLM biases. Fig. 4 and 14 demonstrate that LLM outputs tend to allocate more attention to text tokens. Even for original image-text pairs where the number of image tokens exceeds that of text tokens, the sum of attention scores for text surpasses . An intriguing observation is that when we feed the original text-image pair (Naive), pure text (None), and fully noisy images (Noise), the model exhibits similar attention patterns. Most attention is directed towards special but uninformative tokens, rather than vision tokens or text sequences, such as the question itself. In open-ended generation tasks, we randomly select two questions from LLaVA-Bench (as shown in Fig. 11) and examine the attention scores as more tokens were generated. As depicted in Fig. 5, with the increasing length of generated text during open-ended tasks, there was a corresponding reduction in attention allocated to images, exacerbating the issue. Consequently, the model becomes increasingly prone to generating content independently of input images, shedding light on the genesis of hallucinated content from an attention perspective. Additionally, the results also indicate that the shallow layers assign higher attention to visual tokens; however, as the feature progresses into deeper layers, attention on vision tokens will be reduced.
@@ 2.2 Model Calibration
Post-hoc Debiasing of Prediction Results (Fig. 3 (left)). In addressing the bias issue, we implement a “calibration” step for the model’s output probabilities using an affine transformation [36, 14, 52]: . In this equation, a weight matrix and a bias vector are applied to the original probabilities, producing adjusted probabilities. For classification tasks, represents the probabilities associated with each label name, normalized to sum up to one. To enhance efficiency, we constrain the matrix to be diagonal, utilizing vector scaling [14, 52]. An intuitive approach to learning is to initially estimate the LLM’s bias towards certain answers by inputting a meaningless vision input , such as None or Noise. The resulting output prediction is denoted as . Ideally, LLMs should assign a uniform distribution score to this test input. For instance, when presented with the question “Is the color of the fox in the image brown?”, supplying a meaningless image or no image should result in equal probabilities for ’yes’ and ’no’. This is because the model cannot make a decision. However, due to the model’s biases, it tends to assign a higher score to ’yes’. This error can be corrected by setting and to an all-zero vector. Subsequently, the debiased result is obtained as . In our experimental findings, we establish that employing unrelated vision inputs (Noise, Zero, One) does not consistently enhance model performance and exhibits sensitivity to question types. Consequently, as a default debiasing approach, we opt for image-free methods, specifically None and Unk. Additionally, we explore a debiasing method using Both image-free inputs. In this configuration, we set to None and , obtaining respective probability distributions and . We then calculate the probability distribution .
@@
Visual Debias Decoding (VDD) (Fig. 3 (Right)). While post-hoc debias methods can be naturally extended to open-ended generation settings with the entire vocabulary as the label space, implementation faces challenges due to the large label space. At the same time, it’s crucial to maintain the correlation or co-occurrence of output text tokens and simply enforcing a uniform LM output distribution could significantly harm generation quality. In response, we adopt the concept of contrastive decoding [25, 21, 50, 42] and introduce the VDD strategy. Similar to post-hoc debiasing, we feed the original image and text into the LVLM to obtain logits and feed image-free input to get logits , where contains only text priors. To emphasize the contribution of visual information, we aim to mitigate undesired behaviors highlighted by the image-free logits and generate text based on the remaining positive behaviors of the LVLM when provided with the image input. To implement this concept, we propose the contrastive objective: . In this context, represents the amplification level of the debiasing method. It is crucial to recognize that the logits from pure-text models are not invariably incorrect. They can adeptly capture various fundamental aspects of English grammar and common sense. Therefore, adopting a universal penalty for all tokens may not be suitable in every instance. To tackle this nuanced challenge, we introduce an adaptive plausibility constraint across the output vocabulary of LVLMs. This constraint is intricately linked to the confidence level associated with the output distribution when original visual inputs are considered: Here, is a hyperparameter that truncates the next token distribution of . A larger entails more aggressive truncation, keeping only high probability tokens. During generation, we set the logits of all tokens not in to :
@@ 2.3 Impact of Decoding Configuration on Model Performance
We observe that existing open-source LVLMs undergo evaluations using a variety of generation strategies, as outlined in Tab. 5. This variability in evaluation configuration introduces a potential source of performance variability, impacting the thorough exploration of LVLM capabilities. Our research primarily delves into decoding configurations, specifically experimenting with distinct strategies:
@@ 3.2 Systematic Evaluation
Impact of Post-hoc Debias and Debias Sampling Methods on LVLM Performance: We conduct a comprehensive analysis by combining various debiasing strategies, including both post-hoc debias methods and debias sampling strategies, to investigate their influence on LVLM truthfulness and reasoning abilities. The results obtained from the MME dataset are presented in Tab. 1. Our key observations can be summarized as follows: (1) Post-hoc debias methods significantly enhance model truthfulness by mitigating hallucination, with both None and Unk strategies exhibiting superior performance compared to the naive approach, particularly in mitigating object and attribute hallucination; (2) Debias sampling strategies strike a better balance, with VDD-None and VDD-Unk achieving competitive hallucination scores while consistently outperforming in reasoning scores; (3) Combining different post-hoc debiasing strategies (Both) does not yield significant benefits, whereas combining debias sampling strategies shows promising results; (4) Post-hoc debiasing methods operate independently from sampling methods such as VCD and our proposed debias sampling methods. For instance, VCD (u), which combines VCD sampling with Unk post-hoc debiasing, outperforms VCD alone. Notably, VCD exhibits subpar performance with the LLaVA-13B model, likely due to its sensitivity to image noise levels. In contrast, our model relies on text rather than images, resulting in both higher efficiency and robustness. Similar patterns can also be observed in other datasets such as POPE (see Fig. 6), where the proposed debiasing methods and sampling techniques consistently outperform the naive and other baseline methods. Additionally, we notice that employing image-based debiasing methods such as Noise/Zero does not consistently yield improvements. In our experiments, we observe a much larger variance and sensitivity to question type with such methods, further emphasizing the robustness of our image-free debiasing strategies.
@@
Effect of Decoding Strategies on LVLM Performance: In Tab. 2 (left), we conduct a comprehensive assessment of the LLaVA-v1.5 model’s performance across three distinct datasets—POPE-MSCOCO (average F1 across three subsets), MME (total score across 3 evaluation subsets), and MMMU (average ACC across 36 subsets)—by employing various decoding strategies and settings. The “Default” column denotes the baseline performance using default sampling parameters, while the “Sampling” columns exhibit results obtained through diverse decoding strategies, including temperature (Temp ), Top-, and Top-. The “Overall” column highlights the optimal performance achieved across all sampling strategies on all subsets. Notably, the application of the optimal sampling strategy leads to a substantial improvement in the overall model performance compared to default configurations. To underscore the generality of this phenomenon beyond LLaVA models, we extend our evaluation to include Qwen-VL and InstructBLIP in the POPE-MSCOCO benchmark. The results consistently demonstrate performance enhancements with refined sampling strategies across existing LVLMs. In Tab. 2 (right), we emphasize the superior performance of our proposed method, labeled as “Ours”, in comparison to alternative strategies. Our method surpasses fine-tuned models on specific instruction datasets, highlighting the effectiveness of our refined sampling strategies. Moreover, the 7B model, employing these refined sampling strategies, outperforms the 13B model with different fine-tuning strategies, including both supervised fine-tuning and RLHF. Additional detailed results for all subsets can be found in the appendix (Tabs. 6, 8, 9 and 10). Furthermore, a detailed comparison with additional baseline models on the POPE three subsets is provided in Tab. 11.
| Setting | Default | Sampling | Overall | |||
| Dataset | LLaVA-v1.5 | Temp | Top- | Top- | ||
| POPE MSCOCO | 7B | 79.5 | 84.0 | 84.0 | 84.0 | 84.1 |
| 13B | 77.2 | 84.2 | 84.2 | 84.1 | 84.2 | |
| Qwen-VL | 81.8 | 83.6 | 83.7 | 83.3 | 83.7 | |
| InstructBLIP | 78.9 | 79.8 | 84.5 | 79.9 | 84.5 | |
| MMMU | 7B | 34.4 | 43.4 | 40.3 | 40.8 | 44.7 |
| 13B | 36.3 | 43.2 | 43.4 | 43.1 | 45.8 | |
| MME-OH | 7B | 313.3 | 363.3 | 345.0 | 355.0 | 363.3 |
| 13B | 340.0 | 355.0 | 340.0 | 355.0 | 355.0 | |
| MME-AH | 7B | 260.0 | 323.3 | 301.7 | 325.0 | 330.0 |
| 13B | 303.3 | 321.7 | 311.7 | 321.7 | 321.7 | |
| MME-RE | 7B | 317.9 | 409.6 | 443.2 | 423.6 | 449.6 |
| 13B | 295.4 | 428.9 | 405.7 | 415.0 | 453.9 | |
@@
Overall Performance: Upon reaching the final iteration of our proposed method, we initially identify the optimal sampling configuration for all backbones. Subsequently, we apply debias sampling methods. Simultaneously, for classification tasks, we implement our post-hoc debiasing method, utilizing both None and Unk inputs concurrently. As depicted in Fig. 7, it becomes evident that VCD fails to consistently improve all baseline models, particularly in reasoning tasks. In contrast, our proposed method demonstrates superior results across all aspects and backbone models. Finally, the proposed method is assessed on the LLava-Bench, and the results presented in Fig. 8 demonstrate that, for open-ended generation tasks, VCD does not surpass the default decoding configuration consistently, although it occasionally outperforms in the complex subset. In contrast, our proposed VDDconsistently exhibits improvements over the default decoding strategy. The right columns of the VDD section illustrate that, for generation tasks, model performance is consistently influenced by various decoding configurations. Unlike multi-choice QA tasks, which favor a low temperature (), we observe that generation tasks benefit from a higher or a larger value. Detailed results and results with larger backbones are shown in Tabs. 16 and 17. Qualitative examples in Sec. 0.C.1 demonstrate that VDD is more helpful and less hallucinated, generating more precise illustrations.

### v2

# https://github.com/yfzhang114/LLaVA-Align Debiasing Multimodal Large Language Models

@@ Abstract
LLMs such as GPT-4 [34], PaLM [7], and LLaMA-2 [47] have evolved to handle both text and images. This expansion into the multimodal domain involves further pre-training with image-text pairs or fine-tuning with specialized vision instruction tuning datasets, resulting in the birth of powerful Large Visual Language Models (LVLMs). These models excel at understanding complex visual patterns and translating them into coherent language. However, despite their capabilities, our investigation reveals a notable issue. The content generated by LVLMs is significantly biased towards the underlying LLMs used during pre-training, rather than being influenced by the input images. Even in situations where the image is entirely noisy or absent, LVLMs confidently generate answers, indicating a bias towards the learned language patterns. This observation is illustrated in Fig.˜1, where querying an LVLM about the color of a non-existent fox produces confident yet unrealistic answers like brown, gray, or blue. To explore this bias further, we conduct a toy experiment using questions about color, shape, number, and relationships for 80 entities in the MSCOCO dataset. Our experimental baselines include Noise (fully noisy image and question input), None (no image provided), and Zeros/Ones (image with all zero/one values). Fig.˜1 displays the top 15 choices and corresponding probabilities of frequently occurring answers across these cases. Remarkably, even when no image is available or questions are about non-existent or meaningless images, current LVLMs tend to produce specific answers. This indicates a substantial bias originating from the pretraining on LLMs. Such biases bring persistent challenges, particularly the issue of hallucination [13, 26, 26, 43], pose a significant concern, impacting the reliability and applicability of LVLMs. In this paper, we also aim to uncover the underlying reasons behind such a phenomenon, and the results will provide insights into the design of current LVLMs and their training strategies.
@@
In this paper, we present two training-free strategies to mitigate generation biases associated with underlying LLMs. Our first approach introduces a “calibration” step to adjust the output distribution. Specifically, we assess the model’s inclination toward specific answers by inputting a dummy test with no image, as illustrated in Fig.˜1, where we replace the image with ’None’ or meaningless tensors. We then determine the calibration parameters to ensure that the image-free input produces uniform scores for each answer without yielding overly confident outputs. This calibration procedure establishes suitable parameters without requiring additional training data, referred to as the ‘‘Post-Hoc debias’’ approach, proving intuitive for classification tasks111In binary classification settings like POPE, we aim for the input with a meaningless image to have ’yes’ and ’no’ with an equal probability of .. However, in open-ended generation, calibrating each token’s logits uniformly, even without image input, is unreasonable due to the inherent correlation between text tokens. In such cases, we extend the Post-Hoc debias method to its ‘‘Debias sampling’’ counterpart. Inspired by contrastive decoding methods [25, 21], we compute the difference between generation token log-probabilities for one with a correct image and another with a meaningless vision input. Both the Post-Hoc debias and debias sampling methods serve as regularization techniques, reducing the generation results’ dependency on pure text or meaningless image input.
@@
Moreover, we have identified a significant level of instability in the performance of current LVLMs across various generative configurations. Our primary hypothesis is that existing evaluations are predominantly based on a default decoding setting, limiting the exploration of the model’s full capabilities. As extensively illustrated in Fig.˜2, diverse choices in generative configurations yield substantially different performance outcomes. Notably, adjusting the generation temperature can lead to a notable increase in the average F1 score of POPE-MSCOCO, shifting from 76 to approximately 84.04. This improvement surpasses the previous results achieved by the 13B model and the RLHF-tuned model by a considerable margin, indicating that the current estimation of LVLM capability may be underestimated. Furthermore, our investigation reveals that different models exhibit preferences for distinct generation configurations, highlighting the instability in the current evaluation of LVLMs. This raises concerns about the fairness of evaluations, particularly since they often rely on default generation configurations or selectively choose the best configuration for the proposed model. To maximize the full potential of existing LVLMs, we systematically investigate six LVLM models across four benchmarks, meticulously searching their generative configurations to identify optimal settings. Our results significantly outperform previously reported ones by a substantial margin, underscoring the importance of selecting the best decoding configuration for each model.
@@ 2 Method
Background of Large Vision-Language Models. A LVLM with parameters , conditioned on a query text , a vision input , parametrizes a probability distribution over response sequences . Each response token is sampled from a token probability distribution, namely , where denotes the logits of . Usually, an LLM is first pre-trained on a large, unlabeled text dataset, then an LVLM utilizes the LLM as the text encoder and incorporates vision representations in different ways [3, 23, 9, 12]. Related work about LVLMs and Hallucination in LVLMs can be found in Appendix˜0.A.
@@ 2.1 Exploring the Impact of LLM Biases on LVLMs
To assess the influence of LLM biases on LVLMs, we conduct toy experiments with five target questions detailed in Tab.˜4. For each entity of the MSCOCO dataset (appendix Tab.˜3), we design corresponding prompts and pair them with different types of vision inputs, including unrelated vision input such as fully noisy images (Noise) or entirely black or white images (Zero/One), or replace vision input with pure text, where we either remove all vision tokens (None) or replace them with a meaningless placeholder like </unk> (Unk). As illustrated in Fig.˜1 and in the appendix Fig.˜13, LVLMs exhibit biases towards specific answers, such as associating “white” with color-related questions, “round” with shape-related questions, and “many” with number-based questions. We further analyze attention mechanisms to explain why models heavily rely on LLM biases. Fig.˜4 and 14 demonstrate that LLM outputs tend to allocate more attention to text tokens. Even for original image-text pairs where the number of image tokens exceeds that of text tokens, the sum of attention scores for text surpasses . An intriguing observation is that when we feed the original text-image pair (Naive), pure text (None), and fully noisy images (Noise), the model exhibits similar attention patterns. Most attention is directed towards special but uninformative tokens, rather than vision tokens or text sequences, such as the question itself. In open-ended generation tasks, we randomly select two questions from LLaVA-Bench (as shown in Fig.˜11) and examine the attention scores as more tokens were generated. As depicted in Fig.˜5, with the increasing length of generated text during open-ended tasks, there was a corresponding reduction in attention allocated to images, exacerbating the issue. Consequently, the model becomes increasingly prone to generating content independently of input images, shedding light on the genesis of hallucinated content from an attention perspective. Additionally, the results also indicate that the shallow layers assign higher attention to visual tokens; however, as the feature progresses into deeper layers, attention on vision tokens will be reduced.
@@ 2.2 Model Calibration
Post-hoc Debiasing of Prediction Results (Fig.˜3 (left)). In addressing the bias issue, we implement a “calibration” step for the model’s output probabilities using an affine transformation [36, 14, 52]: . In this equation, a weight matrix and a bias vector are applied to the original probabilities, producing adjusted probabilities. For classification tasks, represents the probabilities associated with each label name, normalized to sum up to one. To enhance efficiency, we constrain the matrix to be diagonal, utilizing vector scaling [14, 52]. An intuitive approach to learning is to initially estimate the LLM’s bias towards certain answers by inputting a meaningless vision input , such as None or Noise. The resulting output prediction is denoted as . Ideally, LLMs should assign a uniform distribution score to this test input. For instance, when presented with the question “Is the color of the fox in the image brown?”, supplying a meaningless image or no image should result in equal probabilities for ’yes’ and ’no’. This is because the model cannot make a decision. However, due to the model’s biases, it tends to assign a higher score to ’yes’. This error can be corrected by setting and to an all-zero vector. Subsequently, the debiased result is obtained as . In our experimental findings, we establish that employing unrelated vision inputs (Noise, Zero, One) does not consistently enhance model performance and exhibits sensitivity to question types. Consequently, as a default debiasing approach, we opt for image-free methods, specifically None and Unk. Additionally, we explore a debiasing method using Both image-free inputs. In this configuration, we set to None and , obtaining respective probability distributions and . We then calculate the probability distribution .
@@
Visual Debias Decoding (VDD) (Fig.˜3 (Right)). While post-hoc debias methods can be naturally extended to open-ended generation settings with the entire vocabulary as the label space, implementation faces challenges due to the large label space. At the same time, it’s crucial to maintain the correlation or co-occurrence of output text tokens and simply enforcing a uniform LM output distribution could significantly harm generation quality. In response, we adopt the concept of contrastive decoding [25, 21, 50, 42] and introduce the VDD strategy. Similar to post-hoc debiasing, we feed the original image and text into the LVLM to obtain logits and feed image-free input to get logits , where contains only text priors. To emphasize the contribution of visual information, we aim to mitigate undesired behaviors highlighted by the image-free logits and generate text based on the remaining positive behaviors of the LVLM when provided with the image input. To implement this concept, we propose the contrastive objective: . In this context, represents the amplification level of the debiasing method. It is crucial to recognize that the logits from pure-text models are not invariably incorrect. They can adeptly capture various fundamental aspects of English grammar and common sense. Therefore, adopting a universal penalty for all tokens may not be suitable in every instance. To tackle this nuanced challenge, we introduce an adaptive plausibility constraint across the output vocabulary of LVLMs. This constraint is intricately linked to the confidence level associated with the output distribution when original visual inputs are considered: Here, is a hyperparameter that truncates the next token distribution of . A larger entails more aggressive truncation, keeping only high probability tokens. During generation, we set the logits of all tokens not in to :
@@ 2.3 Impact of Decoding Configuration on Model Performance
We observe that existing open-source LVLMs undergo evaluations using a variety of generation strategies, as outlined in Tab.˜5. This variability in evaluation configuration introduces a potential source of performance variability, impacting the thorough exploration of LVLM capabilities. Our research primarily delves into decoding configurations, specifically experimenting with distinct strategies:
@@ 3.2 Systematic Evaluation
Impact of Post-hoc Debias and Debias Sampling Methods on LVLM Performance: We conduct a comprehensive analysis by combining various debiasing strategies, including both post-hoc debias methods and debias sampling strategies, to investigate their influence on LVLM truthfulness and reasoning abilities. The results obtained from the MME dataset are presented in Tab.˜1. Our key observations can be summarized as follows: (1) Post-hoc debias methods significantly enhance model truthfulness by mitigating hallucination, with both None and Unk strategies exhibiting superior performance compared to the naive approach, particularly in mitigating object and attribute hallucination; (2) Debias sampling strategies strike a better balance, with VDD-None and VDD-Unk achieving competitive hallucination scores while consistently outperforming in reasoning scores; (3) Combining different post-hoc debiasing strategies (Both) does not yield significant benefits, whereas combining debias sampling strategies shows promising results; (4) Post-hoc debiasing methods operate independently from sampling methods such as VCD and our proposed debias sampling methods. For instance, VCD (u), which combines VCD sampling with Unk post-hoc debiasing, outperforms VCD alone. Notably, VCD exhibits subpar performance with the LLaVA-13B model, likely due to its sensitivity to image noise levels. In contrast, our model relies on text rather than images, resulting in both higher efficiency and robustness. Similar patterns can also be observed in other datasets such as POPE (see Fig.˜6), where the proposed debiasing methods and sampling techniques consistently outperform the naive and other baseline methods. Additionally, we notice that employing image-based debiasing methods such as Noise/Zero does not consistently yield improvements. In our experiments, we observe a much larger variance and sensitivity to question type with such methods, further emphasizing the robustness of our image-free debiasing strategies.
@@
Effect of Decoding Strategies on LVLM Performance: In Tab.˜2 (left), we conduct a comprehensive assessment of the LLaVA-v1.5 model’s performance across three distinct datasets—POPE-MSCOCO (average F1 across three subsets), MME (total score across 3 evaluation subsets), and MMMU (average ACC across 36 subsets)—by employing various decoding strategies and settings. The “Default” column denotes the baseline performance using default sampling parameters, while the “Sampling” columns exhibit results obtained through diverse decoding strategies, including temperature (Temp ), Top-, and Top-. The “Overall” column highlights the optimal performance achieved across all sampling strategies on all subsets. Notably, the application of the optimal sampling strategy leads to a substantial improvement in the overall model performance compared to default configurations. To underscore the generality of this phenomenon beyond LLaVA models, we extend our evaluation to include Qwen-VL and InstructBLIP in the POPE-MSCOCO benchmark. The results consistently demonstrate performance enhancements with refined sampling strategies across existing LVLMs. In Tab.˜2 (right), we emphasize the superior performance of our proposed method, labeled as “Ours”, in comparison to alternative strategies. Our method surpasses fine-tuned models on specific instruction datasets, highlighting the effectiveness of our refined sampling strategies. Moreover, the 7B model, employing these refined sampling strategies, outperforms the 13B model with different fine-tuning strategies, including both supervised fine-tuning and RLHF. Additional detailed results for all subsets can be found in the appendix (Tabs.˜6, 8, 9 and 10). Furthermore, a detailed comparison with additional baseline models on the POPE three subsets is provided in Tab.˜11.
| Setting | Default | Sampling | Overall | |||||
| Dataset | LLaVA-v1.5 | Temp | Top- | Top- | ||||
7B | 79.5 | 84.0 | 84.0 | 84.0 | 84.1 | ||
| 13B | 77.2 | 84.2 | 84.2 | 84.1 | 84.2 | |||
| Qwen-VL | 81.8 | 83.6 | 83.7 | 83.3 | 83.7 | |||
| InstructBLIP | 78.9 | 79.8 | 84.5 | 79.9 | 84.5 | |||
| MMMU | 7B | 34.4 | 43.4 | 40.3 | 40.8 | 44.7 | ||
| 13B | 36.3 | 43.2 | 43.4 | 43.1 | 45.8 | |||
| MME-OH | 7B | 313.3 | 363.3 | 345.0 | 355.0 | 363.3 | ||
| 13B | 340.0 | 355.0 | 340.0 | 355.0 | 355.0 | |||
| MME-AH | 7B | 260.0 | 323.3 | 301.7 | 325.0 | 330.0 | ||
| 13B | 303.3 | 321.7 | 311.7 | 321.7 | 321.7 | |||
| MME-RE | 7B | 317.9 | 409.6 | 443.2 | 423.6 | 449.6 | ||
| 13B | 295.4 | 428.9 | 405.7 | 415.0 | 453.9 | |||
@@
Overall Performance: Upon reaching the final iteration of our proposed method, we initially identify the optimal sampling configuration for all backbones. Subsequently, we apply debias sampling methods. Simultaneously, for classification tasks, we implement our post-hoc debiasing method, utilizing both None and Unk inputs concurrently. As depicted in Fig.˜7, it becomes evident that VCD fails to consistently improve all baseline models, particularly in reasoning tasks. In contrast, our proposed method demonstrates superior results across all aspects and backbone models. Finally, the proposed method is assessed on the LLava-Bench, and the results presented in Fig.˜8 demonstrate that, for open-ended generation tasks, VCD does not surpass the default decoding configuration consistently, although it occasionally outperforms in the complex subset. In contrast, our proposed VDDconsistently exhibits improvements over the default decoding strategy. The right columns of the VDD section illustrate that, for generation tasks, model performance is consistently influenced by various decoding configurations. Unlike multi-choice QA tasks, which favor a low temperature (), we observe that generation tasks benefit from a higher or a larger value. Detailed results and results with larger backbones are shown in Tabs.˜16 and 17. Qualitative examples in Sec.˜0.C.1 demonstrate that VDD is more helpful and less hallucinated, generating more precise illustrations.